OpenStudy (anonymous):

Find the common difference for the sequence shown 1/4, 5/16, 3/8, ... 1/16 1/4 1/2

OpenStudy (michele_laino):

the requested difference \(d\), is: \[\Large \begin{gathered} r = \frac{5}{{16}} - \frac{1}{4} = ...? \hfill \\ \hfill \\ r = \frac{3}{8} - \frac{5}{{16}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

oops... \[\Large \begin{gathered} d = \frac{5}{{16}} - \frac{1}{4} = ...? \hfill \\ \hfill \\ d = \frac{3}{8} - \frac{5}{{16}} = ...? \hfill \\ \end{gathered} \]